OpenStudy (anonymous):

Divide: (12+9i)/(3-4i)

OpenStudy (anonymous):

\[\frac{ 12+9i }{ 3-4i }\times \frac{ 3+4i }{ 3+4i }=\frac{ (12+9i)(3+4i) }{ 9+16 }\]

OpenStudy (anonymous):

Rationalise the denominator mate....

OpenStudy (anonymous):

Thanks forgot about that. U cant have i in the denominator. i was using long division. fail.

OpenStudy (anonymous):

Mate. If you need the answer to refer to when you do it. Here's what you want to get as your answer. 3i is the answer.

OpenStudy (anonymous):

I knew what the answer was. I just forgot how to get it.

OpenStudy (anonymous):

ah okay, no worries then.

OpenStudy (anonymous):

Can you help me with one more problem?

OpenStudy (anonymous):

How do i find an absolute value of a binomial with an i?

OpenStudy (anonymous):

absolute value of a complex number means the distance of the curve of the complex number. To find the distance of complex number curve, you just use pythagoras... \[\left| a+ib \right|=\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

So "i" just disappears?

OpenStudy (anonymous):

Dude haven;t you learnt about absolute values? The absolute value of something is the distance from the origin.

OpenStudy (anonymous):

You're finding the distance from the origin to to the point (a, b)

OpenStudy (anonymous):

@blahblah_who_cares